38. Does Christine now reject or fail to reject (FTR) Ho and for what reason? a) FTR because Its<tel So is the study inconclusive2 The class decides that the sample size was t0o small and believe many more joeys should be sampled Well Christine cant do this all by herself (finding all these ioeys and weighing them is ghaustingworki soshe and2 of herclassmates all goto ustralia and eacho themfin sandweighs 23. jovs think about that- what is the sample size going to be?) They compile all their data and determine the combined average weight to be 504.09 mg with the sample s d still equal to 40 mg They perform an hvpothesis test using aOL 39. What test do they use and why? a) t-test because its a sample s.d 40. The tests not significant at α 01 However, using the p-value, how confident are we ofthe test result? a) 95.24% When Europeans first came to Australia, they thought that Koalas looked like bears and in fact that they were bears so they were often called Koala bears. Koalas are not bears-they are marsupials and their correct name is Koalas Though often called the koala bear, this cuddly animal is not a bear at all, it is a marsupial, or pouched mammal. After giving birth, a female koala carries her baby in her pouch for about six months. www.todayiloundout com
Answer #1

38) Since |ts| < |tc|

If |ts| > |tc| we reject the hypothesis but |ts| < |tc| we don't have evidence to reject the hypothesis.

39)

It is given the sample data and we have given sample standard deviation so t test is used

40)

C = 1 - 0.01 = 0.99

> qt(.99,23)
[1] 2.499867
> 2.4999*40/sqrt(23)
[1] 20.85061
> 20.85/504
[1] 0.04136905

1 - 0.04136905 = 0.9586

Hence it is above 95%